# SciAssess: Benchmarking LLM Proficiency in Scientific Literature Analysis

### v3

@@ C.9 Compound Disease Recognition
Example Paragraph Twenty children with acute lymphoblastic leukemia who developed meningeal disease were treated with a high-dose intravenous methotrexate regimen that was designed to achieve and maintain CSF methotrexate concentrations of 10(-5) mol/L without the need for concomitant intrathecal dosing. The methotrexate was administered as a loading dose of 6,000 mg/m2 for a period of one hour followed by an infusion of 1,200 mg/m2/h for 23 hours. Leucovorin rescue was initiated 12 hours after the end of the infusion with a loading dose of 200 mg/m2 followed by 12 mg/m2 every three hours for six doses and then every six hours until the plasma methotrexate level decreased to less than 1 X 10(-7) mol/L. The mean steady-state plasma and CSF methotrexate concentrations achieved were 1.1 X 10(-3) mol/L and 3.6 X 10(-5) mol/L, respectively. All 20 patients responded to this regimen, 16/20 (80%) achieved a complete remission, and 20% obtained a partial remission. The most common toxicities encountered were transient serum transaminase and bilirubin elevations, neutropenia, and mucositis. One patient had focal seizures and transient hemiparesis but recovered completely. High-dose intravenous methotrexate is an effective treatment for the induction of remission after meningeal relapse in acute lymphoblastic leukemia.
We prompt the model with the following:
@@
Prompt System Message: You are an expert in the field of Biomedical. You are a biologist AI. I’ll give you the abstract of literature. Please identify all the (compound,disease) relations in the abstract, and just give me a list of all relations you recognized. To be mentioned, all the relations must be strictly presented to me in the format ’(compound 1, disease 1),(compound 2, disease 2),(compound 3, disease 3),….’, without adding any additional comments or explanations! User Message: [processed text] Expected Answer: (methotrexate, transient hemiparesis), (methotrexate, neutropenia), (methotrexate, seizures), (methotrexate, mucositis)
C.10 Gene Disease Function
Examples of process text:
@@ C.10 Gene Disease Function
Prompt System Message: You are an expert in the field of Biomedical. In this semantic role recognition task, you need to follow 3 steps, and finally just return me triples that needed. First, you need to identify the entities in the text. Entities can be classified into 2 categories–molecular, and trigger word. ’Molecular’ includes disease, gene, protein, and enzyme. ’Trigger word’ includes 1)Variation (Var), which means DNA, RNA, and mutations in proteins and changes in molecular structure, e.g. ’mutations on the Arg248 and Arg282’, ’mutant R282W’, ’missense mutations’; 2)Molecular Physiological Activity (MPA), including molecular activity, gene expression and molecular physiological activity, e.g. ’phosphorylation’, ’transcription’, ’histone methylation’, ’bioactivation of cyclophosphamide’; 3)Interaction, molecule-to-molecule or molecule-to-cell connections, e.g. ’bind’, ’interaction’; 4)Pathway, e.g. ’Bmp pathway’,’PI3K pathway’; 5)Cell Physiological Activity (CPA), Activities at or above the cellular level, including cellular reactivity and cell or organ development and growth, e.g. ’T helper cell responses’, ’renal development’; 6)Regulation (Reg), a neutral cue word or phrase meaning no loss or gain, e.g. ’resolved in’, ’regulated’; 7)Positive Regulation (PosReg), a cue word or phrase that indicates the acquisition of a function, e.g. ’facilitates’, ’enhanced’, ’increased’; 8)Negative Regulation (NegReg), a clue word or phrase that indicates a loss of function, e.g. ’suppressed’, ’decreased’, ’inhibited’. Second, you need to identify the semantic role labeling objects, including ’ThemeOf’(from the main thing entity to the current entity) and ’CauseOf’(From the current entity to the Cause entity). Third, please give me triples that contain entities and semantic role labeling objects(ThemeOf or Causeof). User Message: [processed text] Expected Answer: (frameshift, CauseOf, reduction), (caused by, CauseOf, lower), (mutation, CauseOf, results in), (beta(0) thalassaemia, ThemeOf, results in), (beta globin mRNA, ThemeOf, reduction), (beta0 thalassaemia gene, ThemeOf, frameshift), (insertion, CauseOf, generates), (premature chain terminator, ThemeOf, generates), (amounts of beta globin mRNA, ThemeOf, lower), (mutation, CauseOf, caused by), (degradation, ThemeOf, caused by).
The ideal answer are "(frameshift,CauseOf,reduction),(caused by,CauseOf,lower),(mutation,CauseOf,results in),(beta(0) thalassaemia,ThemeOf,results in),(beta globin mRNA,ThemeOf,reduction),(beta0 thalassaemia gene,ThemeOf,frameshift),(insertion,CauseOf,generates),(premature chain terminator,ThemeOf,generates),(amounts of beta globin mRNA,ThemeOf,lower),(mutation,CauseOf,caused by),(degradation,ThemeOf,caused by)".
C.11 Gene Disease Regulation
Examples of process text:
Example Paragraph Mutation spectra in autosomal dominant and recessive retinitis pigmentosa in northern Sweden. Retinal degenerations represent a heterogeneous group of disorders affecting the function of the retina. The frequency of retinitis pigmentosa (RP) is 1/3500 worldwide, however, in northern Sweden it is 1/2000 due to limited migration and a ’founder’ effect. In this study we identified genetic mechanisms underlying autosomal dominant and recessive RP present in northern Sweden. Several novel mutations unique for this region were found. In an autosomal recessive form of RP, Bothnia dystrophy caused by mutations in the RLBP1 gene, bi-allelic mutations R234W, M226K and compound heterozygosity, M226K+R234W was detected.In dominant form of RP mapped to 19q13.42 a 59 kb genomic deletion including the PRPF31 and three other genes was found.These data provide additional information on the molecular mechanisms of RP evolvement and in the future might be useful in development of therapeutic strategies. Identification of the disease-causing mutations allowed introducing molecular genetic testing of the patients and their families into the clinical practice.
For extracting triplets (entities, regulatory types, entities), we prompt the model with:
Prompt System Message: You are an expert in the field of Biomedical. In this Gene-Disease relation extraction task, you need to follow 3 steps. You need to extract the (gene, function change, disease) triplet from the text, such as: (SHROOM3, LOF, Neural tube defects). The second element in the triple means the regulation that the gene produces to the disease. Types of regulations are: LOF and GOF, which indicate loss or gain of function; REG, which indicates a general regulatory relationship; COM, which indicates that the functional change between genes and diseases is more complex, and it is difficult to determine whether the functional change is LOF or GOF. Please return all the relations extracted from the text in ternary format (GENE, FUNCTION, DISEASE). If there are more than one triple, please write in this form: ’(GENE, FUNCTION, DISEASE),(GENE, FUNCTION, DISEASE)’ User Message: [processed text] Expected Answer: (RLBP1, REG, Bothnia dystrophy).
C.12 Affinity Extraction
An example output is shown in Table 7.
Table 7: Example output of affinity data extraction task
Compound Name SMILES Affinities
2 / C1[C@H](O[C@H]([C@H]1F) N2C=CC(=NC2=O)N)CO >200000 nM 4000 nM
We prompt the model with the following:
Prompt System Message: You are an expert in the field of pharmaceutical chemistry, and your task is to summarize the results of activity assays from an article in a tabular format. Please follow these steps to complete the task: 1. Determine if the article includes an activity assay. If it does, locate the section(s) presenting the assay results, which are usually in one or more tables. 2. Compile all the activity assay results into a single table. You may use multiple columns to represent different conditions or outcomes of various experiments. 3. Identify the names or codes used in the table, such as Example 1 or Compound A, and find the corresponding sections in the article that mention these substances. Extract the full name and SMILES notation of each substance. 4. Compile the names and SMILES notations of each substance in the table. Output in CSV format with multi-index (Affinities, protein/cell line), write units not in the header but in the value like "10.5 µM". Quote the value if it has comma! For example: “‘csv Compound,Name,SMILES,Affinities,Affinities,Affinities,Affinities ,,,5HT1A (IC50),5HT1D (IC50),5HT-UT (IC50),5HT1E (<affinity type>) "5a","Aspirin","CC(=O)Oc1ccccc1C(=O)O",2.0 nM,8.0 nM,12.6 nM,>1000 nM “‘ 5. If there are multiple tables, concat them. Don’t give me reference or using "…", give me complete table!
C.13 Tag to Molecule
Prompt System Message: You are an expert in the field of organic chemistry, can help user get SMILES formula from documents. Give me only the answer without explanation. If you can’t answer the completion, just reply "Unknown". User Message: What’s the SMILES formula of molecule "Sumatriptan"? Expected Answer: "CNS(=O)(=O)CC1=CC2=C(C=C1)NC=C2CCN(C)C"
C.14 Markush to Molecule
Prompt System Message: You are an expert in the field of chemistry, can help user insert substituents into CXSMILES-type markush formula to get SMILES formula (removing Hs), just reply SMILES without explanation. User Message: *C(*)CC(*)CC* |$A;;Pol_{p};;;Q_{e};;;M_{p}$|, A = H, Pol = NH2, Q = OH, M = [Li]
C.15 Molecule in Document
Prompt System Message: You are an expert in the field of chemistry. You are given a SMILES formula of a molecule, and should judge whether it is in the document. If the molecules are given by Markush formula (containing R group), You need to 1) analyze the skeletons of the provided molecule and the molecule in the literature or patent, and 2) if the compare the variable values of the molecular structure with the range of variable values given in the met, to determine whether the molecule is covered by the literature or patent. Answer in "Yes" or "No". User Message: [ document.pdf ] Does the molecule "CC(CCCCCCCC1=CC(=C(C(=C1)OC)OC)OC)CCC(C2=CC=CS2)O" appear in the document?
C.16 Reaction QA
Prompt System Message: You are an expert in the field of organic chemistry. User Message: Which compound is in the reactants or reagents of the following reaction? a) O=C(C(C)C(OC)=O)C1CC1 b) c4ccc(B3OB(c1ccccc1)OB(c2ccccc2)O3)cc4 c) CC(=O)OP(=O)([O-])[O-].[NH4+].[NH4+] d) Nc1[nH]c(F)nc2ncnc12 Only write the options and values down, such as ’b) 2045.1’. The new reaction you should deal with is the Synthesis of (R)-2-ethynyl glyceraldehyde 3-phosphate (5) using co-immobilized PanK-Rd4BB and AcK kinases. Expected Answer: c) CC(=O)OP(=O)([O-])[O-].[NH4+].[NH4+]
C.17 Target Extraction
Prompt System Message: You are an expert in the field of biochemistry, please show which targets cell lines this paper’s experiments do against. Here is an output demonstration: ’A549, HepG2’ User Message: The file content relating to the query is as follows: [ docoment.pdf ] Expected Answer: MV4-11, MOLM-13, THP-1, U937, HL-60
C.18 Electrolyte Table QA
We prompt the model with the following:
@@ C.18 Electrolyte Table QA
Prompt System Message: You are an expert in the electrolytes field. Please answer the following multiple choice question correctly. Only write the options and values down, such as ’(b) 4.301’. User Message: In the upper paper, what are the minimum and maximum intramolecular distances (nm) of dimethyl carbonate? a) 0.41/0.87 b) 0.49/0.67 c) 0.25/0.25 d) 0.25/0.38 Expected Answer: a) 0.41/0.87
C.19 OLED Property Extraction
An example output is shown in Table 8.
@@ C.19 OLED Property Extraction
Prompt Please give a complete list of Host, Host’s SMILES structure (if exists), Dopant, Assistant Dopant (if exists), Td/Tg/ET, Von,max EQE/CE/PE,EQE/CE/PE, and CIE [x, y] * Output in csv format with columns of those attributes, do not write units only the value like "10.5". * Quote the column name or Host’s Name or Dopant’s Name if it contains space or special characters like ",". * If there are multiple tables, concat them. Don’t give me reference or using "…", give me complete table! * Should return all columns mentioned, if empty just return ‘NaN‘. "Host" and "Dopant" should not be empty. * "Host" and "Dopant" should be short name of the organic molecule. * Should find more information from the whole content, including tables, text. for example, you should return: “‘csv Host,SMILES,Dopant,Td [°C]/Tg [°C]/ET [eV], Von [V],max EQE [%]/CE [cd A-1]/PE [lm W-1],EQE [%]/CE [cd A-1]/PE [lm W-1],"CIE [x, y]" PPO1,O=P(c1ccccc1)(c1ccccc1)c1ccccc1,FCNIr,–/74/3.02,–,17.1/20.5/14.3,–/–/–,"(0.14, 0.16)" PPO2,O=P(c1ccccc1)(c1ccccc1)c1ccccc1,FCNIr,–/123/3.02,–,18.4/21.1/16.6,–/–/–,"(0.14, 0.15)" “‘
C.20 Polymer Composition Extraction
@@
Prompt System Message: You are an expert in the field of polymer solar cells researcher who answers the following multiple choice question correctly. Only write the options and values down, such as ’b) 2045.1’. User Message: In this paper, What is the blend ratio of donor to acceptor in the most efficient solar cell? a) 1:4 b) 20:8 c) 30:50 d) 2:4
C.21 Polymer Property Extraction
@@
Prompt User Message: You are an expert in the field of polymer solar cells researcher. Please give a complete list of Nickname, $\text{PCE}_{\text{max}}$(%), $\text{PCE}_{\text{ave}}$(%), Voc (V), Jsc (mA $\text{cm}^{2}$), FF; * Output in csv format with columns of those attribution, do not write units only the value like "10.5". * If there are multiple tables, concat them. Don’t give me reference or using "…", give me complete table! * Should return all columns mentioned, if empty just return ‘NaN‘. Nickname should not be empty. * Nickname should be short name of polymers, for example: ‘PCBM:PffBT4T-2OD:PC61PM‘ should return ‘PffBT4T-2OD‘. * Only return acceptor ‘PC71BM‘ related records. * If with different experiment settings for the same nickname, only return the record with ‘highest PCE‘ ! * Should find more information from the whole content, including tables, text. * For FF use 0.xx instead of xx.x, for example: 63.0 should return 0.63 ! for example, you should return: “‘csv Nickname, $\text{PCE}_{\text{max}}$(%), $\text{PCE}_{\text{ave}}$(%), Voc (V), Jsc (mA $\text{cm}^{2}$), FF PBTTT-C14,2.34,2.34,0.53,9.37,0.48 “‘
C.22 Solubility Extraction

### v5

@@ C.1.4 Compound Disease Recognition
Example Paragraph Twenty children with acute lymphoblastic leukemia who developed meningeal disease were treated with a high-dose intravenous methotrexate regimen that was designed to achieve and maintain CSF methotrexate concentrations of 10(-5) mol/L without the need for concomitant intrathecal dosing. The methotrexate was administered as a loading dose of 6,000 mg/m2 for a period of one hour followed by an infusion of 1,200 mg/m2/h for 23 hours. Leucovorin rescue was initiated 12 hours after the end of the infusion with a loading dose of 200 mg/m2 followed by 12 mg/m2 every three hours for six doses and then every six hours until the plasma methotrexate level decreased to less than 1 X 10(-7) mol/L. The mean steady-state plasma and CSF methotrexate concentrations achieved were 1.1 X 10(-3) mol/L and 3.6 X 10(-5) mol/L, respectively. <rest of the paragraph>.
We then prompt the model with the following:
@@
Prompt System Message: You are a biologist AI. I’ll give you the abstract of literature. Please use chain of thought reasoning to identify all the (compound, disease) relations in the abstract. First, analyze the abstract step by step, explaining your reasoning for identifying each relation. Then, provide a final list of the relations in the format: ’(compound 1, disease 1),(compound 2, disease 2),(compound 3, disease 3),….’, without adding any additional comments or explanations. Format: Reasoning: [Reasoning] Answer: [List of recognized relations] User Message: [processed text] Expected Answer: (methotrexate, transient hemiparesis), (methotrexate, neutropenia), (methotrexate, seizures), (methotrexate, mucositis)
C.1.5 Disease Entities Recognition
Similarly, this task involves recognizing disease entity names using data from Wei et al. (2016) and additional expert-annotated data, evaluating the performance of LLMs in identifying specialized disease names:
Prompt System Message: You are an expert in the field of Biomedical. You are a biologist AI. I’ll give you the abstract of literature. Please use chain of thought reasoning to identify all the disease entities in the abstract. First, analyze the abstract step by step, explaining your reasoning for identifying each disease entity. Then, provide a final list of the disease entities you recognized in the format: (disease 1), (disease 2), (disease 3). Format: Reasoning: [Reasoning] Answer: [List of recognized diseases] User Message: In unanesthetized, spontaneously hypertensive rats the decrease in blood pressure and heart rate produced by intravenous clonidine, 5 to 20 micrograms/kg, was inhibited or reversed by nalozone, 0.2 to 2 mg/kg. The hypotensive effect of 100 mg/kg alpha-methyldopa was also partially reversed by naloxone. Naloxone alone did not affect either blood pressure or heart rate. In brain membranes from spontaneously hypertensive rats clonidine, 10(-8) to 10(-5) M, did not influence stereoselective binding of [3H]-naloxone (8 nM), and naloxone, 10(-8) to 10(-4) M, did not influence clonidine-suppressible binding of [3H]-dihydroergocryptine (1 nM). These findings indicate that in spontaneously hypertensive rats the effects of central alpha-adrenoceptor stimulation involve activation of opiate receptors. <rest of the input>. Expected Answer: (hypertensive), (hypotensive)
C.1.6 Gene Disease Function
The Gene Disease Text Mining task focuses on "Gene-Disease" association semantics text mining. It evaluates the ability of models to extract and understand relationships between genes and diseases from scientific literature, with a focus on identifying gene and disease entities Ouyang et al. (2022). Examples of process text:
@@ C.1.6 Gene Disease Function
Prompt System Message: You are an expert in the field of Biomedical. In this semantic role regconition task, you need to follow 3 steps, and finally just return me triples that needed. First, you need to identify the entities in the text. Entities can be classified into 2 categories–molecular, and trigger word. ’Molecular’ includes disease, gene, protein, and enzyme. ’Trigger word’ includes: 1)Variation(Var), which means DNA, RNA, and mutations in proteins and changes in molecular structure, e.g. ’mutations on the Arg248 and Arg282’, ’mutant R282W’, ’missense mutations’; 2)Molecular Physiological Activity (MPA), including molecular activity, gene expression and molecular physiological activity, e.g. ’phosphorylation’, ’transcription’, ’histone methylation’, ’bioactivation of cyclophosphamide’; 3)Interaction, molecule-to-molecule or molecule-to-cell connections, e.g. ’bind’, ’interaction’; 4)Pathway, e.g. ’Bmp pathway’,’PI3K pathway’; 5)Cell Physiological Activity (CPA), Activities at or above the cellular level, including cellular reactivity and cell or organ development and growth, e.g. ’T helper cell responses’, ’renal development’; 6)Regulation (Reg), a neutral cue word or phrase meaning no loss or gain, e.g. ’resolved in’, ’regulated’; 7)Positive Regulation (PosReg), a cue word or phrase that indicates the acquisition of a function, e.g. ’facilitates’, ’enhanced’, ’increased’; 8)Negative Regulation (NegReg), a clue word or phrase that indicates a loss of function, e.g. ’suppressed’, ’decreased’, ’inhibited’. Second, you need to identify the semantic role labeling objects, including ’ThemeOf’(from the main thing entity to the current entity) and ’CauseOf’(From the current entity to the Cause entity). Third, please give me tripples that contain entities and semantic role labeling objects(ThemeOf or Causeof). Use chain of thought reasoning to explain your process of identifying the entities and relations, and then provide the final triples in the format: (…), (…) Format: Reasoning: [Reasoning] Answer: [List of recognized triples] User Message: [processed text] Expected Answer: (frameshift, CauseOf, reduction), (caused by, CauseOf, lower), (mutation, CauseOf, results in), (beta(0) thalassaemia, ThemeOf, results in), (beta globin mRNA, ThemeOf, reduction), (beta0 thalassaemia gene, ThemeOf, frameshift), (insertion, CauseOf, generates), (premature chain terminator, ThemeOf, generates), (amounts of beta globin mRNA, ThemeOf, lower), (mutation, CauseOf, caused by), (degradation, ThemeOf, caused by).
C.2 Chemistry
C.2.1 MMLU-Pro-Chemistry
The example of MMLU-Pro-Chemistry is similar to the MMLU-Pro-Biology task in Appendix C.1.1.
C.2.2 Electrolyte Table QA
The composition and properties of organic electrolytes are crucial for battery performance, stability, and safety. To evaluate the model’s retrieval capabilities regarding electrolyte information, we posed multiple-choice questions about the components of solution systems and the dissolution reactions, focusing on their physical and chemical properties as presented in the tables within the articles. We prompt the model with the following:
@@ C.2.2 Electrolyte Table QA
Prompt System Message: You are an expert in the electrolytes field. Please answer the following multiple choice question correctly. Use chain of thought reasoning and provide reasoning before selecting the correct answer (e.g., a)xxx, or b)xxx). Format: Reasoning: [Reasoning] Answer: [Answer] User Message: In the upper paper, what are the minimum and maximum intramolecular distances (nm) of dimethyl carbonate? a) 0.41/0.87 b) 0.49/0.67 c) 0.25/0.25 d) 0.25/0.38 Expected Answer: a) 0.41/0.87
C.2.3 OLED Property Extraction
This task evaluates the LLM’s ability to extract information about OLED molecules and their optical properties. It tests several key capabilities, including their understanding of complex and domain-specific language and their ability to interpret and extract data from tables. An example output is shown in Table 5.
@@ C.2.3 OLED Property Extraction
Prompt System Message: You are an expert in the field of organic photovoltics. Please give a complete list of Host, Host’s SMILES structure (if exists), Dopant, Assistant Dopant (if exists), Td/Tg/ET, Von,max EQE/CE/PE,EQE/CE/PE, and CIE [x, y] * Output in csv format with columns of those attributes, do not write units only the value like "10.5". * Quote the column name or Host’s Name or Dopant’s Name if it contains space or special characters like ",". * If there are multiple tables, concat them. Don’t give me reference or using "…", give me complete table! * Should return all columns mentioned, if empty just return ‘NaN‘. "Host" and "Dopant" should not be empty. * "Host" and "Dopant" should be short name of the organic molecule. * Should find more information from the whole content, including tables, text. for example, you should return: “‘csv Host,SMILES,Dopant,Td /Tg /ET, Von,max EQE/CE/PE,EQE/CE/PE,"CIE" PPO1,O=P(c1ccccc1)(c1ccccc1)c1ccccc1,FCNIr,–/74/3.02,–, 17.1/20.5/14.3,–/–/–,"(0.14, 0.16)" PPO2,O=P(c1ccccc1)(c1ccccc1)c1ccccc1,FCNIr,–/123/3.02,–, 18.4/21.1/16.6,–/–/–,"(0.14, 0.15)" “‘ Please use a step-by-step approach to analyze the content and ensure that all relevant information is accurately extracted. Only provide reasoning for how you identified each attribute and output the final csv format. Format: Reasoning: [Reasoning] Answer: [Extracted csv] User Message: [ docoment.pdf ]
C.2.4 Polymer Chart QA
The processing steps and properties of polymer materials are often represented through charts. Extracting information from these charts and integrating it with textual data is crucial. To further assess the retrieval capabilities of models concerning polymer chart information, we designed multiple-choice questions involving polymer composition, processing techniques, and properties.
The example of Polymer Chart QA can be found in a similar format to the Biology Chart QA task in Appendix C.1.2.
C.2.5 Polymer Composition QA
This task involves extracting the blend ratio of donor to acceptor in the most efficient solar cell from the text of scientific literature.
@@ C.2.5 Polymer Composition QA
Prompt System Message: You are an expert in the field of polymer solar cells researcher who answers the following multiple choice question correctly. Use chain of thought reasoning and provide reasoning before selecting the correct answer (e.g., a)xxx, or b)xxx). Format: Reasoning: [Reasoning] Answer: [Answer] User Message: In this paper, What is the blend ratio of donor to acceptor in the most efficient solar cell? a) 1:4 b) 20:8 c) 30:50 d) 2:4 Expected Answer: a) 1:4
C.2.6 Polymer Property Extraction
This task focuses on extracting vital values such as power conversion efficiency (PCE) and open-circuit voltage ($V_{\mathrm{OC}}$) from tables within the literature.
@@ C.2.6 Polymer Property Extraction
Prompt System Message: You are an expert in the field of polymer solar cells researcher. Please give a complete list of Nickname, PCE_max, PCE_ave, Voc , Jsc, FF; * Output in csv format with columns of those attribution, do not write units only the value like "10.5". * If there are multiple tables, concat them. Don’t give me reference or using "…", give me complete table! * Should return all columns mentioned, if empty just return ‘NaN‘. Nickname should not be empty. * Nickname should be short name of polymers, for example: ‘PCBM:PffBT4T-2OD:PC61PM‘ should return ‘PffBT4T-2OD‘. * Only return acceptor ‘PC71BM‘ related records. * If with different experiment settings for the same nickname, only return the record with ‘highest PCE‘ ! * Should find more information from the whole content, including tables, text. * For FF use 0.xx instead of xx.x, for example: 63.0 should return 0.63 ! for example, you should return: “‘csv Nickname,PCE_max(%),PCE_ave(%),Voc (V),Jsc (mA $cm^{2}$),FF PBTTT-C14,2.34,2.34,0.53,9.37,0.48 “‘ Please use a step-by-step approach to analyze the content and ensure that all relevant information is accurately extracted. Only provide reasoning for how you identified each attribute and output the final csv format. Format: Reasoning: [Reasoning] Answer: [Extracted csv] User Message: [ docoment.pdf ]
C.2.7 Solubility Extraction
Organic electrolytes, extensively used in battery technologies, comprise organic solvents, lithium salts, and additives. Understanding solubility in organic electrolytes is crucial as it impacts the efficiency of electrolytic processes, product selectivity, and equipment design. This task evaluates the LLM’s capability in retrieving solubility-related tables. Papers typically select data from various aspects to describe the system, making it challenging to combine multiple tables for fuzzy matching. Therefore, we focus on examining the LLM’s semantic understanding ability, enabling the model to select the most relevant and comprehensive table related to “solubility” from numerous alternatives and convert it into the specified format.